OpenStudy (anonymous):

PLEASE HELPPPPPPPPPPPPPP !!!!!!!!!!!!!! what is the solution to x=sqrt-x+6 a. 2 & -3 b. -2 & 3 c. -3 d. 2

OpenStudy (kc_kennylau):

you actually mean sqrt-x as in \(\sqrt{-x}\)?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

oh wait you actually mean \(\sqrt{-x+6}\) not \(\sqrt{-x}+6\)

OpenStudy (kc_kennylau):

Square both sides :)

OpenStudy (anonymous):

on the question its the first one

OpenStudy (kc_kennylau):

Just square both sides first :) can we do this step by step?

OpenStudy (anonymous):

To be honest I have NOOOOOOO idea what im doing ! Im failing algebra 2 EXTREMELY !!! but if you can go through it thats fine too.. ILL TRY LOL :)

OpenStudy (kc_kennylau):

Could you square both sides? :)

OpenStudy (anonymous):

would the square root be -x ?

OpenStudy (kc_kennylau):

well the sqrt isn't just to -x but to (-x+6) as a whole...

OpenStudy (kc_kennylau):

So we have: \[x=\sqrt{-x+6}\] Squaring both sides and we get: \[x^2=-x+6\]

OpenStudy (kc_kennylau):

any problem till here? :)

OpenStudy (anonymous):

i dont understand how you get the x ?

OpenStudy (kc_kennylau):

wait are you in a hurry

OpenStudy (anonymous):

lol, well im taking care of my little baby, thats why im so slow on replying ....

OpenStudy (kc_kennylau):

i'm also helping other ppl too, so i may also be slow :)

OpenStudy (anonymous):

ok, so how do i get the x ?

OpenStudy (kc_kennylau):

Do you know how to do quadratic equations? :)

OpenStudy (anonymous):

I SUCK AT THOSE

OpenStudy (kc_kennylau):

http://www.youtube.com/watch?v=2ZzuZvz33X0 and http://www.youtube.com/watch?v=bNQY0z76M5A will help you :D (sorry for the length of these videos)

OpenStudy (kc_kennylau):

I'm not providing the direct answer, because the answer is useless. What you need is not the answer, it's how to do more questions of this kind.

OpenStudy (anonymous):

i went through each one & cant find it

OpenStudy (anonymous):

x=2, so its D

OpenStudy (anonymous):

and im 100% sure

OpenStudy (kc_kennylau):

\[x=\sqrt{-x+6}\]\[x^2=-x+6\]\[x^2+x-6=0\]\[(x-2)(x+3)=0\]\(x-2=0\) or \(x+3=0\) \(x=2\) or \(x=-3\) Therefore it's A and i'm 100% sure.

OpenStudy (anonymous):

Lol ! wow . 2 different answers and both of u are sure , what to do lol

OpenStudy (kc_kennylau):

I'm 200% sure so believe me LOL

OpenStudy (kc_kennylau):

nah just kidding the check is here: http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=x%3Dsqrt(-x%2B6)

OpenStudy (kc_kennylau):

LOL HE WINS sorry my maths sucks